What is the slope of the line through (-7,-8)(−7,−8)left parenthesis, minus, 7, comma, minus, 8, right parenthesis and (0,4)(0,4)

Answer by math_helper(2461)

slope = rise / run = (change in y value) / (change in x value) = (y2-y1)/(x2-x1)
You are free to choose which point to assign to (x1,y1):
Choosing (x1,y1) = (-7,-8) ( and therefore (x2,y2)=(0,4) ):
slope = (4-(-8)) / (0- (-7)) = 12 / 7 or about 1.7143 to 4 decimal places.

You can reverse the assignments of (x1,y1) and (x2,y2) to prove to yourself that the slope calculation is unaffected.
Slope is usually represented by the letter 'm' in the equations of lines.